OpenStudy (anonymous):

If customer A has an appointment 2 2/3 hours before customer B who has an appointment 5/6 of an hour before customer C. How much later is customer C‘s appointment than customer A’s? 2 1/2 hours Not Here 2 hours 3 hours 3 1/2 hours

OpenStudy (anonymous):

2 2/3 = 8/3 so; (8/3 ) + (5/6) = 21/6 >> 3 hours